A 50-kg sack of rice, resting on the ground, is dragged by a 100-N force at 30 degrees with the horizontal. What is the acceleration of the sack?


1
Expert's answer
2021-03-08T08:10:55-0500

I assume there is no friction in this problem.

The horizontal component of the "F = 100N" force is:


"F_x = F\cos\theta"

where "\theta = 30\degree."

According to the second Newton's law, the acceleration provided by the net force acting on a body (in horizontal direction) is:


"a = \dfrac{F_x}{m} = \dfrac{F\cos\theta}{m}"

where "m = 50kg" is the mass of the sack. Thus, obtain:


"a =\dfrac{100\cos30\degree}{50} = \sqrt{3}\space m/s^2\approx 1.73\space m/s^2"

Answer. "\sqrt{3}\space m/s^2\space \text{or}\space 1.73\space m/s^2"